### v1

# Anomaly detection with flow-based fast calorimeter simulators

Abstract
Recently, several normalizing flow-based deep generative models have been proposed to accelerate the simulation of calorimeter showers. Using CaloFlow as an example, we show that these models can simultaneously perform unsupervised anomaly detection with no additional training cost. As a demonstration, we consider electromagnetic showers initiated by one (background) or multiple (signal) photons. The CaloFlow model is designed to generate single photon showers, but it also provides access to the shower likelihood. We use this likelihood as an anomaly score and study the showers tagged as being unlikely. As expected, the tagger struggles when the signal photons are nearly collinear, but is otherwise effective. This approach is complementary to a supervised classifier trained on only specific signal models using the same low-level calorimeter inputs. While the supervised classifier is also highly effective at unseen signal models, the unsupervised method is more sensitive in certain regions and thus we expect that the ultimate performance will require a combination of approaches.
I Introduction
In 2012, the final piece of the Standard Model, the Higgs boson, was discovered at the Large Hadron Collider (LHC) by the ATLAS [1] and CMS [2] collaborations. Despite this milestone, compelling theoretical and experimental reasons continue to drive the search for physics Beyond the Standard Model (BSM). Regrettably, the extensive search programs conducted by ATLAS [3, 4, 5], CMS [6, 7, 8], and LHCb [9] at the LHC have yet to yield conclusive evidence for new BSM physics. Given the impossibility of conducting dedicated searches for every conceivable BSM scenario, most LHC searches target specific signals derived from theoretical priors, leaving substantial portions of the LHC phase space unexplored. This limitation has spurred the development of more model-agnostic search strategies, with the hope of detecting BSM physics at the LHC. The advent of deep learning has given rise to various model-agnostic anomaly detection methods designed to explore uncharted territories of the LHC phase space — for reviews and original references, see e.g. [10, 11, 12, 13, 14].
Normalizing flows [15, 16, 17, 18, 19, 20] represent a class of deep learning models particularly valuable for generative modeling and density estimation tasks. A normalizing flow is characterized by a parametric diffeomorphism $f_{\theta}$ mapping between a latent space, with a known distribution $\pi(z)$, and a data space of interest with an analytically unknown distribution $p(x)$. In the context of a conditional normalizing flow, this transformation becomes $f_{\theta}(x|c)$, where $c$ denotes the conditional inputs to the flow. It is defined through a series of invertible functions, parameterized by $\theta$, that can be trained by maximizing the log-likelihood of the data following the change of variables formula:
$\log\bigl(p(x|c)\bigr{missing})=\log(\pi(f_{\theta}(x|c)))+\log|\det\left(\mathcal{J}\bigl{(}f_{\theta}(x|c)\bigr{)}\right)\Bigr{|},$
where $\mathcal{J}\bigl{(}f_{\theta}(x|c)\bigr{)}$ represents the Jacobian of the transformation $f_{\theta}(x|c)$. The allowable transformations must also have a computationally tractable Jacobian, ideally efficient to compute, and the probability density of the base distribution $\pi(z)$ must be known. A common choice for $\pi(z)$ is the standard normal distribution.
Recently, normalizing flows have found successful applications in fast calorimeter simulation tasks [39, 40, 43, 42, 44, 51]. In this paper, we demonstrate the utility of these flow-based generative models as unsupervised anomaly detectors for identifying BSM physics in calorimeter shower data. Specifically, we apply CaloFlow [39], a flow-based fast calorimeter simulation model, to single-photon showers from a new sampling calorimeter version [52] of the CaloGAN dataset [53, 21, 22]. In this context, single-photon showers serve as the background events, while the signal events consist of photon showers originating from a generic BSM particle $\chi$ that undergoes the decay $\chi\to\gamma\gamma$. The $\chi$ particle is taken to be invisible and only interacts indirectly with the calorimeter through its decay products. By training CaloFlow to maximize the log-likelihood when evaluated on background events, we are able to detect the signal events based on a cut on the log-likelihood. We focus on achieving signal sensitivity, but the approach could be combined with a number of background estimation strategies [54, 11].
While we believe this is the first unification of simulation and anomaly detection, both subjects have been well-studied with machine learning. Many deep generative models have been studied for calorimeter simulation [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 35, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50], including a number of proposals developed on the CaloChallenge datasets [55], and they are also now being integrated into experimental workflows [36]. We focus on normalizing flows since they give direct access to the likelihood. This information can also be extracted from diffusion models [56] and it would be interesting to compare approaches in the future. For anomaly detection, unsupervised methods have been extensively studied (e.g. Refs. [57, 58, 59] and many others) and also include density-based approaches [12]. Like the density-based methods, we use the likelihood directly as the anomaly score.111 Note that this is not unique and is sensitive to how the data are represented/preprocessed [60, 61, 62].
There may be other ways of reusing the generative model for BSM searches, including fine-tuning supervised models based on particular signal hypotheses.
This paper is organized as follows. In Sec. II, we describe calorimeter setup and the datasets that were used during training/evaluation. In Sec. III, we explain how CaloFlow is used as an unsupervised anomaly detector by placing cuts on the log-likelihood of background and signal showers evaluated using CaloFlow. In Sec. IV, we include the results of performing unsupervised anomaly detection with CaloFlow. Finally, we conclude in Sec. V.
II Dataset
For this study, we decided to create a new, more realistic sampling calorimeter version [52] of the CaloGAN dataset. The original dataset included energy contributions from both active and inactive calorimeter layers, whereas our new dataset only includes energy contributions from the active layers as would be available in practice. The sampling fraction of our new calorimeter setup is $\sim 20\%$. The simple detector is a three-layer, liquid argon (LAr) sampling calorimeter cube with 480mm side-length that is inspired by the ATLAS LAr electromagnetic calorimeter [63]. It is simulated as follows: Geant4 [64, 65, 66] is used to generate particles and simulate their interaction with our calorimeter using the Ftfp_Bert physics list based on the Fritiof [67, 68, 69, 70] and Bertini intra-nuclear cascade [71, 72, 73] models with the standard electromagnetic physics package [74]. While we use this new simulator to create a dataset for anomaly detection, we expect it should be more generally useful for a broad variety of tasks.
The calorimeter showers are represented as three-dimensional images that are binned in position space. In this representation, the calorimeter shower geometry is made up of voxels (volumetric pixels) and the details of the calorimeter voxel dimensions are included in Table 1. Figure 1 (taken from [21, 75]) shows the three-dimensional representation of a shower in the CaloGAN calorimeter. The three longitudinal layers are separated in the figure for visualization purposes. In this work, the center of the detector is taken to be at $z=0$ m, while the front face of layer 0 is positioned at $z=1$ m which is consistent with an ATLAS-like configuration.
Layer
index
$z$ length
$\eta$ length
$\phi$ length
Number
of voxels
0 90 5 160 $3\times 96$
1 347 40 40 $12\times 12$
2 43 80 40 $12\times 6$
Table 1: Dimensions of a calorimeter voxel. The positive $z-$axis (radial direction in full detector) is the direction of particle propagation, the $\eta$ direction is along the proton beam axis, and $\phi$ is perpendicular to $z$ and $\eta$. For the number of voxels, the first (second) number is the number of bins in the $\phi$ ($\eta$) direction (e.g., $12\times 6$ refers to 12 $\phi$ bins and 6 $\eta$ bins).
Figure 1: Three-dimensional representation of the shower in the CaloGAN calorimeter, figure taken from  [21, 75]. Not-to-scale separation among the longitudinal layers is added for visualization purposes.
For the background dataset (single-photon calorimeter showers), we generate 100k showers with incident energies $E_{\text{inc}}$ uniformly distributed in the range $[1,100]$ GeV. This dataset was used as the training dataset for CaloFlow where we used a train/validation split of 70%/30%. A second independent dataset of 100k showers with the same range of $E_{\text{inc}}$ was generated and used for evaluation.
To obtain the signal datasets, we defined the hypothetical $\chi$ particle in Geant4 which has the same properties as the $\eta^{0}$ particle except having a different mass and being invisible to the detector. Next, we generate multiple sets of 100k showers that originate from $\chi\to\gamma\gamma$ decay at 10 chosen fixed displacements from the center of detector along the $z-$axis. The chosen fixed displacements are $z_{i}\in\{0.33,0.66,1.00,1.04,1.08,1.16,1.24,1.32,1.40,1.44\}$ m. Note that the first two displacements are located in front of the calorimeter, while the last eight displacements are located at distinct positions within the calorimeter. The energy of the $\chi$ particle was fixed at 50 GeV. Such a hypothetical scenario might arise from the decay of a 100 GeV particle (close to e.g. the Higgs boson mass), which is at rest, to a pair of $\chi$ particles.
[…]
To consider particles with fixed lifetimes, we construct new signal datasets using the fixed displacement samples. In particular, we use the probability that the $\chi$ particle decays at various positions along the $z-$axis to determine the proportion of showers originating from decays at each of the 10 fixed displacements. Generating directly with fixed lifetimes would have been too inefficient, since we are discarding decays after the calorimeter volume, which can happen often for these lifetimes that we consider. More details of the fixed lifetime signal datasets are included in Section IV.2.
We also generated 100k signal showers with kinetic energy [^2] We distinguish between kinetic energy and incident energy for massive particles. The incident energy is the sum of the kinetic energy and rest mass energy of the particle. This distinction only becomes noticeable in the case of $m_{\chi}=5$ GeV. uniformly distributed in the range [1,100] GeV for a $\chi$ particle with $m_{\chi}=5\times 10^{-3}$ GeV and lifetime $\tau=1.00$ ns. A second set of 100k signal showers was generated with the same range of kinetic energies, but for a $\chi$ particle with $m_{\chi}=5$ GeV and lifetime $\tau=1.00$ ns. Each of these two datasets (together with the background dataset) was used to train a supervised classifier described in Section IV.3. Using a range of kinetic energies ensured that we obtain showers with a range of $E_{\text{inc}}$ such that the $E_{\text{inc}}$ would not be used as a discriminating factor by the supervised classifier. This allows for a fairer comparison between performance of the supervised classifier and our unsupervised method since CaloFlow learns the likelihood conditioned on $E_{\text{inc}}$.
III CaloFlow
CaloFlow [39, 40] is an approach to fast calorimeter simulation based on conditional normalizing flows. In the context of fast calorimeter simulation, CaloFlow uses a two-flow method to learn to generate the voxel level shower energies $\vec{\mathcal{I}}$ conditioned on the corresponding incident energies of the showers $E_{\text{inc}}$ denoted by $p(\vec{\mathcal{I}}|E_{\text{inc}})$. Flow-I is constructed to learn the probability density of calorimeter layer energies [^3] The layer energy of a given calorimeter layer is the sum of all the voxel energies in that layer. $E_{i}$ conditioned on the incident energy $p_{1}(E_{0},E_{1},E_{2}|E_{\text{inc}})$, while Flow-II is designed to learn the probability density of the normalized voxel level shower energies conditioned on the calorimeter layer energies $p_{2}\left(\hat{\mathcal{I}}|E_{0},E_{1},E_{2},E_{\text{inc}}\right)$. By normalized, we mean that the voxel energies in each layer are made to sum to unity. The dimensions of the conditional inputs and outputs of the two flows are shown in Table 2. Importantly, these flows were trained only using single photon showers.
One important difference with original CaloFlow is that in this application to anomaly detection we do not have direct empirical access to $E_{\text{inc}}$. For a given shower, we do not know a priori what the corresponding $E_{\text{inc}}$ is and would instead have to use a reconstructed estimate of $E_{\text{inc}}$ which we denote as $E^{(\text{rec})}_{\text{inc}}$. In this work, we use a simple regression method to reconstruct $E_{\text{inc}}$ given the total deposited energy in the calorimeter $E_{\text{dep}}$. The reconstructed incident energy is defined as $E^{(\text{rec})}_{\text{inc}}=\lambda E_{\text{dep}}$, where $\lambda$ is the mean of $E_{\text{inc}}/E_{\text{dep}}$ in the single photon training dataset. Note that the true $E_{\text{inc}}$ is still used for training CaloFlow, while $E^{(\text{rec})}_{\text{inc}}$ is used when performing anomaly detection.
The architecture and training of CaloFlow are outlined in Appendix A. Some modifications were made to the original CaloFlow, while most of the main algorithm remains unchanged.
Conditionals
Dim of
conditional
Output
Dim of
output
Flow-I $E_{\rm inc}$ 1 $E_{0},E_{1},E_{2}$ 3
Flow-II $E_{0},E_{1},E_{2},E_{\rm inc}$ 4 $\hat{\mathcal{I}}$ 504
Table 2: The conditional inputs for each flow, and the features whose probability distributions are the output of each flow.
III.1 Unsupervised anomaly detection with CaloFlow
After training CaloFlow on background single-photon showers, we evaluate the log-likelihood $\log p\left(\hat{\mathcal{I}},E_{i}|E^{(\text{rec})}_{\text{inc}}\right)$ of the background and signal showers by using the trained flows. Using both Flow-I and Flow-II, we are able to obtain the full log-likelihood:
$\log p\left(\hat{\mathcal{I}},E_{i}|E^{(\text{rec})}_{\text{inc}}\right)=\underbrace{\log p_{1}(E_{i}|E^{(\text{rec})}_{\text{inc}})}_{\text{Flow-I}}+\underbrace{\log p_{2}\left(\hat{\mathcal{I}}|E_{i},E^{(\text{rec})}_{\text{inc}}\right)}_{\text{Flow-II}}\,.$ (1)
Figure 2: Plot of log-likelihood comparison between signal showers from $\chi$ decays at the 10 fixed displacements (lines) and background photon showers (filled). This example is for a $\chi$ particle with $m_{\chi}=5$ GeV.
Figure 2 shows an example of a plot of the full log-likelihood of the signal showers from $\chi$ decays at the 10 fixed displacements and background photon showers.
We see from Figure 2 that the signal showers generally have log-likelihoods that are distinguishable from that of the background showers. Hence, we can use CaloFlow as an unsupervised anomaly detector by placing cuts on the full log-likelihood to discriminate signal from background showers. The results are detailed in Section IV.
IV Results
IV.1 Decay at fixed displacement
In this section, we study the effect of varying the displacement from the center of the detector at which the decay occurs on the evaluated likelihood. In reality, particles do not decay at fixed positions, but instead the probability of a particle decaying at a given displacement from where it was created is related to its lifetime $\tau$. Nevertheless, studying the showers produced by the particle $\chi$ at fixed decay positions is interesting from an experimental perspective, and doing so also helps us interpret the more physical results in Section IV.2.
Figure 3: Heatmap of maximum significance improvement $\left(\text{tpr}/\sqrt{\text{fpr}}\right)$ for different masses of the $\chi$ particle $m_{\chi}$ and the 10 fixed displacements where the decay occurs.
Our performance metric for anomaly detection will be the significance improvement which is defined as the signal efficiency (i.e. true positive rate or tpr) divided by the square root of the background efficiency (i.e. false positive rate or fpr). The maximum significance improvement corresponds to the best possible [^4] This is signal model-dependent, but still provides a useful bound on the achievable performance. multiplicative factor by which the signal significance can be improved with a cut on the anomaly score. Figure 3 shows a heatmap of maximum significance improvement for each of the four different particle masses $m_{\chi}$, and the 10 fixed displacements where the decay occurs. In general, we find that showers from decays at larger $z$ are more anomalous. However, there is clear exception in the case of $m_{\chi}=5$ GeV, where the showers originating from decays that occur before the $\chi$ particle reaches the calorimeter (e.g., $z=$ 0.33 m, 0.66 m) are more anomalous than those from decays occuring at the front face of the calorimeter (i.e., $z=1.0$ m). This is due to the fact that the 5 GeV particle is less boosted compared to the other lighter particles that we consider in this study. As a result, the decay of the 5 GeV particle often results in a wider angle between the produced pair of photons which CaloFlow is better able to distinguish from the background single-photon showers. See Figure 4 for an example of such a shower and comparison with a regular photon shower. On the other hand, if a decay occurs right at the front face of the calorimeter, there is insufficient time for the pair of photons to propagate and create two distinct blobs of energy in the calorimeter.
Figure 4: Top row: Example of shower with two distinct blobs of energy originating from a 5 GeV $\chi$ particle. Bottom row: Example of typical photon shower with centralized energy core. The energy deposition in each of the three layers are shown here with layer 0 on the left, layer 1 in the center, and layer 2 on the right.
For the three lowest $m_{\chi}$, we find that CaloFlow struggles to detect signal showers that originate from decays occurring before the front face of the calorimeter (i.e., $z\leq 1.0$ m). These lower mass particles are highly boosted which results in showers from highly collimated photon pairs that are more similar to the background photon showers.
IV.2 Varying lifetime
To assess the performance of our unsupervised anomaly detector on realistic scenarios, we construct new datasets consisting of showers from $\chi$ particles with fixed rest frame lifetimes. In particular, for each chosen mass $m_{\chi}$ and lifetime $\tau$, we have a total of 100k events that are made up of showers from $\chi$ decays at fixed displacements. The proportion of showers associated to decays at each fixed displacement is determined based on the lifetime $\tau$. Particles that decay after the calorimeter volume are not included in the events since they are not detected within the calorimeter.
The probability for a particle to survive for time $t$ before decaying is given by $p_{s}(t)=\exp(-\frac{t}{\gamma\tau})$, where the relativistic boost factor $\gamma=E_{\text{particle}}/M_{\text{particle}}$. Equivalently, the probability that a particle decays before reaching displacement $z$ is given by $p_{d}(z)=1-\exp(-\frac{z}{c\tau\sqrt{\gamma^{2}-1}})$. Since we consider 10 fixed displacements indexed by $i\in\{1,2,...,10\}$, the fraction of the total showers that originates from $\chi$ decay at displacement $z_{i}$ is set to be $\hat{w}_{i}$ which is defined by
$w_{i}=\left\{\begin{array}[]{ll}p_{d}(z_{1}),\quad i=1\\ p_{d}(z_{i})-p_{d}(z_{i-1}),\quad\text{otherwise}\end{array}\right.$
$\text{and}\quad\hat{w}_{i}=\frac{w_{i}}{\sum_{j=1}^{10}w_{j}}$
In other words, the number of showers originating from the decay at $z=z_{i}$ is equal to $10^{5}\times\hat{w}_{i}$ [^5] In some cases, we had to round $\hat{w}_{i}$ to ensure that the total number of showers is equal to 100k.. In this work, we consider three possible lifetimes $\tau\in\{0.01,0.1,1\}$ ns for each choice of $m_{\chi}$.
Figure 5: Plots of significance improvement of our unsupervised (unsup) CaloFlow anomaly detector as a function of background rejection (1/false positive rate) for various types of signal showers originating from $\chi$ particles with different $m_{\chi}$ and $\tau$. Top left: $m_{\chi}=5\times 10^{-3}$ GeV, top right: $m_{\chi}=0.05$ GeV, bottom left: $m_{\chi}=0.5$ GeV, bottom right: $m_{\chi}=5$ GeV. The performance of a random classifier is drawn with black dashed lines to serve as a baseline.
Figure 5 shows the significance improvement of our CaloFlow anomaly detector as a function of the background rejection (1/false positive rate) for various types of signal showers originating from $\chi$ particles with different $m_{\chi}$ and $\tau$.
Overall, we find that CaloFlow is able to achieve significantly better performance compared to a random classifier which we take to be the baseline. One exception is for $(m_{\chi},\tau)=(0.5\ \text{GeV},0.01\ \text{ns})$ where the performance closely matches that of the random baseline. The poorer performance here is due to the lower boost factor $\gamma$ for the $0.5$ GeV particle compared to the two lower particles masses which results in the majority of the decays occurring close to the center of the detector ($z=0$ m). Meanwhile, the mass is still low enough that the photons are not widely separated by the time they reach the calorimeter. As we have seen in Figure 3, such showers are mostly not detected as anomalous by CaloFlow. At longer lifetimes, the showers are more anomalous as a larger proportion of them originate from decays occuring within the calorimeter.
For the three lowest $m_{\chi}$, the significance improvement generally increases with background rejection. Also, we find that significance improvement curves are similar across lifetimes for each of these three lowest $m_{\chi}$. At fixed $m_{\chi}$, the large boost of these particles results in a similar proportion of particles decaying at a given fixed displacement before/within the calorimeter for different lifetimes $\tau$.
The best performance among all the (fixed mass and lifetime) signal models we considered in this study was achieved in the cases with the largest mass of $m_{\chi}=5$ GeV. As explained in Section IV.1, showers from early decay of these more massive particles are quite anomalous according to CaloFlow due to the wider angle between the produced pair of photons. In this case, going to higher lifetimes actually makes these showers slightly less anomalous (which is opposite to the trend seen at lower masses) since it gives the photons less time to separate before interacting with the detector.
IV.3 Comparison with supervised anomaly detection
The previous section showed that the unsupervised anomaly detector has broad coverage across the various model parameters. An important question to ask is how this compares to a dedicated search for the $\chi\rightarrow\gamma\gamma$ signal. For a particular signal model, we would expect that model to outperform the unsupervised approach. However, it is not possible to have a dedicated search for every possible signal and so the key question to ask is how well a supervised model trained on one signal would perform on other signals not seen during training.
In this section, we compare the performance of our method against two supervised classifiers. Each supervised classifier was trained on a combined dataset with 100k signal showers and 100k background showers. The showers originate from particles with kinetic energy uniformly distributed in the range [1,100] GeV. The signal showers originate from $\chi$ particles with lifetime $\tau=1.00$ ns. The first (second) supervised classifier was trained on the a dataset with $m_{\chi}=5\times 10^{-3}\ (5)$ GeV. These models were chosen because they are sufficiently different that they would likely be covered by different dedicated searches. It is thus interesting to ask if a search optimized for one of the models would still be sensitive to the other, since the unsupervised approach has some sensitivity to both models.
The supervised classifier is a fully-connected neural network with two hidden layer with 512 nodes each. We have a 505-dimensional input consisting of the voxel energies normalized by the reconstructed incident energy $\vec{\mathcal{I}}/E^{(\text{rec})}_{\text{inc}}$ (504-dim) and the reconstructed incident energy $E^{(\text{rec})}_{\text{inc}}$ (1-dim). The output layer returns a single number which is passed through a sigmoid activation function. All other activation functions are ReLUs. The supervised classifier was trained for a total of 50 epochs with a train/test/validation split of 60%/20%/20%. These parameters were not extensively optimized, but we found little gain from small variations in the setup.
Figure 6: Comparison plots of significance improvement of our CaloFlow anomaly detector as a function of background rejection (1/false positive rate). The plots for the performance for the $m_{\chi}=5\times 10^{-3}$ GeV and $m_{\chi}=5$ GeV cases are shown on the left and right respectively. Top row: supervised classifier was trained on signal showers originating from $\chi$ particles with $m_{\chi}=5\times 10^{-3}$ GeV and lifetime $\tau=1.00$ ns. Bottom row: supervised classifier was trained on signal showers originating from $\chi$ particles with $m_{\chi}=5$ GeV and lifetime $\tau=1.00$ ns. The performance of our unsupervised CaloFlow anomaly detector (unsup) is shown as solid lines, while that of the supervised classifier (sup) is shown as dotted lines. The performance of a random classifier is drawn with black dashed lines to serve as a baseline.
After training the supervised classifiers, they were evaluated on the signal showers discussed in Section IV.2. Figure 6 shows the comparison plots between the performance of our CaloFlow anomaly detector (unsup) and the supervised classifiers (sup). The top (bottom) row shows the performance of the supervised classifier that was trained on signal showers originating from $\chi$ particles with $m_{\chi}=5\times 10^{-3}\ (5)$ GeV and lifetime $\tau=1.00$ ns.
In the top left plot, we see that the supervised classifier clearly performs better than CaloFlow on the signal showers originating from $\chi$ particles that have same mass ($5\times 10^{-3}$ GeV) as what it was trained on, achieving a larger significance improvement for various background rejection. We found that the supervised classifier still performs well on signal showers originating from $\chi$ particles with higher masses as long as the $\chi$ particles they originate from are still highly boosted.
However, as shown in the top right plot, the performance of the supervised classifier decreases significantly for signal showers associated to $m_{\chi}=5$ GeV that are less highly boosted. We find that CaloFlow achieves a much higher significance improvement for large range of background rejection in the cases of $\tau=0.01,0.1$ ns. For the case of $\tau=1$ ns, the supervised classifier and CaloFlow appear to have comparable performance across different background rejection.
In the bottom right plot, we again see that the supervised classifier outperforms CaloFlow on the signal showers originating from $\chi$ particles that have same mass (5 GeV) as what it was trained on. Not surprisingly, we see that this supervised classifier achieves much higher significance improvement compared to the other supervised classifier that was trained on signal showers associated to $m_{\chi}=5\times 10^{-3}$ GeV (top right plot) when evaluated on signal showers associated to $m_{\chi}=5$ GeV.
Similarly, this supervised classifier evaluated on signal showers associated to $m_{\chi}=5\times 10^{-3}$ GeV (bottom left plot) achieves lower significance improvement compared to the supervised classifier that was trained on signal showers associated to $m_{\chi}=5\times 10^{-3}$ GeV (top left plot). Despite the drop in performance, we see that the supervised classifier still outperforms CaloFlow when evaluated on signal showers it was not trained on. This is likely due to CaloFlow not being able to fully discriminate against signal showers that only decay after the first longitudinal layer (i.e., $E_{0}=0$ GeV), whereas such showers are usually perfectly distinguished by the supervised classifier. As noise is added to the voxel energies when using CaloFlow, this artificially causes $E_{0}>0$ GeV [^6] We checked that a different treatment of the layers with zero energy deposition does not improve the performance. and prevents such showers from being perfectly distinguished by CaloFlow.
This comparison of our unsupervised anomaly detection method against a supervised classifier highlights the potential limitation of model-specific anomaly detection as the supervised model is unable to generalize its excellent performance to signal that is too different from what it was trained on. We note that it is possible to train a supervised classifier on all the types of signal showers we have considered here. Doing so would likely result in the supervised classifier outperforming CaloFlow when evaluated on all the signal showers. However, the point is that even if one is to train a supervised classifier on a large number of signal types, it is impossible to exhaust the space of all possible signals. Hence, there may be an advantage in using model-agnostic, unsupervised anomaly detection methods such as the one we proposed in this work. This is especially true when using flow-based fast calorimeter simulators as no additional training has to be performed to use them as unsupervised anomaly detectors.
V Conclusion and Outlook
Using CaloFlow as an example, we demonstrated how fast calorimeter surrogate models with access to the data likelihood can double up as unsupervised anomaly detectors.
By studying the anomaly detection performance of CaloFlow on showers from $\chi$ particle decays occurring at fixed displacements in the detector, we found that CaloFlow is generally more sensitive to signal showers from decays that occur deeper in the calorimeter. However, in the case of more massive, less highly boosted $\chi$ particles, CaloFlow still has significant discriminative power for showers from decays occurring in front of the calorimeter.
By reweighting the proportion of showers originating from decays at each fixed displacement, we constructed signal datasets corresponding to fixed particle lifetimes. We found that CaloFlow has discriminative power for most of the models we tested. In particular, CaloFlow achieves the best performance in the case with $m_{\chi}=5$ GeV and $\tau=0.01$ ns where the particle is less highly boosted and majority of the particles decay close to the center of the detector.
Finally, we compared to performance of our unsupervised CaloFlow anomaly detector against a supervised classifier. We found that a supervised classifier trained on signal showers from highly boosted $\chi$ particles performed significantly poorer on showers from more massive, less highly boosted particles compared to our unsupervised method. On more boosted signals, the supervised classifier is still highly effective. This highlights the complementarity of different approaches and reaffirms the need to have a diversity of methods in order to achieve broad sensitivity.
Data and code availability
The datasets used in this study can be found at [52] and the software to generate these datasets are located at https://github.com/Ian-Pang/CaloGAN/tree/anomaly_detection. The machine learning software is at https://github.com/Ian-Pang/AD_with_CF.
Acknowledgements
CK would like to thank the Baden-Württemberg-Stiftung for financing through the program Internationale Spitzenforschung, project Uncertainties – Teaching AI its Limits (BWST_IF2020-010). IP and DS are supported by the U.S. Department of Energy (DOE), Office of Science grant DOE-SC0010008 and BN is supported by the DOE under contract DE-AC02-05CH11231.
Appendix A Architecture and training
Here we briefly describe the architecture and training procedure used for CaloFlow (see [39, 40] for more details). There are some differences compared to the implementation in the original CaloFlow papers [39, 40], but most of main algorithm remains the same.
Both Flow-I and Flow-II are Masked Autoregressive Flows (MAFs) [76] with compositions of Rational Quadratic Splines (RQS) [77] as transformations. The RQS transformations are parameterized using neural networks known as MADE blocks [78]. The details of the architecture of Flow-I and Flow-II are summarized in Table 3.
dim of number of layer sizes number of RQS
Model base distribution MADE blocks input hidden output RQS bins tail bound
Flow-I 3 6 64 $2\times 64$ 69 8 14
Flow-II 504 8 378 $1\times 378$ 11592 8 14
Table 3: Summary of architecture of the Flow-I and Flow-II. For the hidden layer sizes, the first number is the number of hidden layers in each MADE block and the second number is the number of nodes in each hidden layer (e.g., $2\times 64$ refers to 2 hidden layers per MADE block with 64 nodes per hidden layer).
The incident energy of the incoming photon is preprocessed as
$E_{\text{inc}}\to\log_{10}(E_{\text{inc}}/10\ \text{GeV})\,.$ (2)
The inputs to the flows are preprocessed as follows:
  • Flow-I: $E_{i}\to 2\left(\log_{10}(E_{i}+1\ \text{keV})-1\right)$
  • Flow-II:
    $E_{i}\to\log_{10}(E_{i}+1\ \text{keV})-2$,
    $\mathcal{I}_{ia}\to u_{\text{logit},ia}(\mathcal{I}_{ia}/E_{i})$,
    where $u_{\text{logit},ia}=\log\frac{\tilde{u}_{ia}}{1-\tilde{u}_{ia}}$,
    $\tilde{u}_{ia}=\alpha+(1-2\alpha)u_{ia}$ and $\alpha=10^{-6}$
The index $i$ denotes the layer number, while the index $a$ specifies the voxel within the given layer. In the original CaloFlow, a different preprocessing was used for the layer energies $E_{i}$ in Flow-I where $E_{i}$ were transformed to unit-space (see [39]).
As in Ref. [39, 40], uniform noise in the range [0,1] keV was applied to the voxel energies during training and evaluation. The addition of noise was found to prevent the flow from fitting unimportant features. The training of both flows in this work is optimized using independent Adam optimizers [79]. Flow-I was trained by minimizing $-\log p(E_{0},E_{1},E_{2}|E_{\text{inc}})$ for 75 epochs with a batch size of 200. Flow-II was trained by minimizing $-\log p\left(\hat{\mathcal{I}}|E_{0},E_{1},E_{2},E_{\text{inc}}\right)$ for 100 epochs with a batch size of 200. The initial learning of $1\times 10^{-4}$ was chosen for the two flows and a multi-step learning schedule was used when training the flows which halves the learning rate after each selected epoch milestone during the training.